### v1

# AdEval: Alignment-based Dynamic Evaluation to Mitigate Data Contamination in Large Language Models

Abstract
As Large Language Models (LLMs) are pretrained on massive-scale corpora, the issue of data contamination has become increasingly severe, leading to potential overestimation of model performance during evaluation. To address this, we propose AdEval (Alignment-based Dynamic Evaluation), a dynamic data evaluation method aimed at mitigating the impact of data contamination on evaluation reliability. AdEval extracts key knowledge points and main ideas to align dynamically generated questions with static data’s core concepts. It also leverages online search to provide detailed explanations of related knowledge points, thereby creating high-quality evaluation samples with robust knowledge support. Furthermore, AdEval incorporates mechanisms to control the number and complexity of questions, enabling dynamic alignment and flexible adjustment. This ensures that the generated questions align with the complexity of static data while supporting varied complexity levels. Based on Bloom’s taxonomy, AdEval conducts a multi-dimensional evaluation of LLMs across six cognitive levels: remembering, understanding, applying, analyzing, evaluating, and creating. Experimental results on multiple datasets demonstrate that AdEval effectively reduces the impact of data contamination on evaluation outcomes, enhancing both the fairness and reliability of the evaluation process.
I Introduction
Large Language Models (LLMs) have significantly advanced the field of Natural Language Processing (NLP) with their remarkable performance across a wide range of tasks [1, 2]. These models rely on pretraining over massive-scale internet corpora, while many widely-used benchmarks are also derived from online resources, inevitably leading to the issue of data contamination [3]. Recent studies indicate that data contamination is prevalent in LLM evaluations [1, 4, 5], undermining the credibility of evaluation results and hindering fair comparisons between models.
To address data contamination, two primary solutions have been proposed: data contamination detection and dynamic data evaluation [6]. Data contamination detection aims to identify overlaps between model outputs and training data [7]. However, these methods face limitations, particularly for proprietary models such as the GPT series, which often incorporate special filtering mechanisms during generation. Additionally, detection methods primarily focus on extreme memorization (i.e., direct reproduction of training data), making it difficult to capture more subtle forms of contamination. Moreover, the training datasets of closed-source models are often treated as trade secrets, leaving external communities with limited ability to intervene directly.
Compared to static benchmarks, dynamic data evaluation effectively circumvents the problem of data contamination, avoiding the inherent limitations of detection methods. Existing dynamic evaluation approaches have achieved notable progress in various domains. For instance, DYVAL [4] introduces dynamic data generation tailored for mathematical tasks; KIEval [6] incorporates multi-turn dialogue-based evaluation; LatestEval [8] excels in generating contamination-free datasets; SciEval [9] emphasizes testing scientific research capabilities through the design of entirely new questions; and [10] propose a practical strategy for generating evaluation data using simple heuristics. These methods explore the potential of dynamic evaluation from different perspectives, offering diverse tools for assessing the capabilities of LLMs.
To address these challenges, this paper proposes AdEval (Alignment-based Dynamic Evaluation), a dynamic data evaluation method designed to reduce the impact of data contamination on LLMs evaluation results by dynamically generating aligned evaluation samples. The key innovations of AdEval are as follows:
  1. 1.
    Dynamically generating aligned evaluation data: AdEval extracts key knowledge points and main ideas from static data and combines them with online search to provide detailed expansions of related content. This process generates dynamic data that aligns with the core concepts of the static data, ensuring both dynamism and consistency.
  2. 2.
    Complexity control and dynamic alignment: AdEval incorporates a complexity regulation mechanism to dynamically adjust the difficulty of questions. This ensures that the generated data matches the difficulty level of static data while accommodating diverse complexity requirements.
  3. 3.
    Multi-dimensional cognitive evaluation: Using Bloom’s taxonomy, AdEval evaluates LLMs across six cognitive dimensions—remembering, understanding, applying, analyzing, evaluating, and creating—offering a comprehensive assessment of their performance in various cognitive tasks.
Experimental results demonstrate that AdEval effectively reduces the impact of data contamination on evaluation outcomes, showcasing significant advantages in the reliability, fairness, and diversity of the generated data.
Figure 1: An Overview of the AdEval Process: Generating Dynamic Datasets from Static Data.
II Related Work
II-A Benchmarks for Evaluating Large Language Models
To evaluate the performance of LLMs across multiple tasks, researchers have proposed various benchmarks. MMLU[11] assesses knowledge and reasoning capabilities across diverse disciplines. HELM[12] expands evaluation dimensions to include aspects such as fairness and efficiency. BIG-Bench[13] introduces 204 challenging tasks to explore the limits of model capabilities. AGIEval[14] and C-Eval[15] focus on standardized examinations and advanced knowledge evaluation in Chinese, respectively.
While these benchmarks cover a wide range of tasks, they primarily rely on static datasets, making them susceptible to data contamination, which can result in an overestimation of model capabilities.
II-B Data Contamination
Data contamination has emerged as a critical challenge in evaluating LLMs[3]. In traditional NLP and ML tasks, separating training and testing data is relatively straightforward. However, as LLMs grow in scale, overlaps between training data and evaluation benchmarks have become increasingly common. This contamination may lead to misunderstandings of model performance and, in some cases, misleading conclusions. Particularly, when evaluation benchmarks and training data are derived from similar online sources, the risk of contamination increases significantly, affecting the fairness and validity of model comparisons.
Several studies have investigated the impact of data contamination. [16] and [17] explored how data contamination during pretraining affects language model performance, analyzing factors such as data volume and model size. [18] and [19] examined the influence of contamination on downstream tasks or zero- and few-shot evaluations. [20] demonstrated a technique called Evasion through Augmented Learning (EAL), an effective contamination method that bypasses most detection strategies.
Recent data contamination detection methods include detecting data leakage through sample or question order manipulation (e.g., PAC[21], [22], [23]); using perplexity-based detection (e.g., [24], [25]); analyzing LLMs’ internal state distributions or identifying low-probability outlier tokens (e.g., DICE [26], Min-K%++[27], [28]); and identifying contamination by detecting and correcting output distributions (e.g., CDD and TED [29]). Other methods, such as DetCon and ComiEval [29], improve benchmark evaluation, while ConStat [30] identifies contamination by comparing benchmark performances. [31] combine guided prompts with classifiers for detection, [32] propose cross-lingual contamination detection, and [33] analyze membership characteristics using statistical inference. Despite their unique contributions, these methods often rely on intrinsic model information or complex analytical techniques, limiting their practical applicability.
In contrast to static benchmarks, dynamic evaluation overcomes data contamination by generating new datasets, providing a more reliable approach for evaluating LLMs. LatestEval [8] and [10] automatically generate test data from up-to-date texts, minimizing overlap with pretraining data. DYVAL [4] and DyVal 2 [34] dynamically produce samples with controllable complexity to evaluate reasoning capabilities and model improvements effectively. SciEval [9] combines static and dynamic data for multi-level evaluation of scientific tasks. [35] employs comprehensive tests to evaluate the subject-specific knowledge of Chinese LLMs.
Additionally, several methods mitigate data contamination and improve evaluation accuracy by generating variants, rephrasing questions, or removing similar samples. For instance, Clean-Eval [36] uses paraphrasing and semantic detection to optimize sample quality. [37] generate variants and rephrase questions for contamination detection, while [38] filters similar samples to reduce overfitting risks. DCQ [39] creates multiple-choice questions to detect data leakage, and KIEval [6] evaluates comprehension and knowledge application through multi-turn dialogues. These studies provide diverse strategies for dynamic evaluation, significantly enhancing fairness and reliability in LLM assessments.
III Methods
The proposed process for dynamically generating datasets is shown in Figure 1 , consisting of the following steps: data preprocessing, extraction of knowledge points and main ideas, online search, question design,and complexity control , quality checking.
III-A Prompt Optimization
Figure 2: Iterative Optimization Process for Prompts
Prompts significantly influence the output quality of large language models (LLMs), as their design directly determines the accuracy and alignment of the generated content with the anticipated answer. To improve output quality, this study adopts a multi-iteration prompt optimization strategy. Utilizing multiple LLMs, such as GPT-4, the prompts are continuously adjusted and refined to guide the model’s outputs closer to the desired results.
As illustrated in Figure 2, the optimization process involves the following key steps:
  1. 1.
    Example Selection: Select 1 to 5 representative examples from a predefined dataset as few-shot examples.
  2. 2.
    Initial Prompt Design: Based on the examples, design the initial version of the prompt, embedding the examples and task requirements into the prompt to form a complete input.
  3. 3.
    Model Generation and Comparison: Input the prompt along with the examples into the LLM to generate answers. Compare the outputs with the standard answers to analyze differences and shortcomings.
  4. 4.
    Optimization Suggestion Generation: Input the initial prompt, the LLM outputs, and the standard answers into another LLM to generate optimization suggestions for the prompt.
  5. 5.
    Prompt Adjustment: Refine and improve the prompt based on the optimization suggestions, guiding the LLM to produce outputs closer to expectations.
III-B Data Preprocessing
Before integrating the data into the AdEval process, it must undergo formatting adjustments and sampling. Considering computational resource and time constraints, uniform sampling is performed on the dataset, selecting 302 samples. Among these, 300 samples are used for dynamically generating new data, while the remaining 2 samples serve as few-shot examples for the prompts.
The unified format of the adjusted dataset is as follows:
    "instruction": "Instruction Content",
    "input": "Input Content",
    "output": "Answer",
    "system": "",
    "history": []
Uniform Sampling Formula:
$i_{k}=\left\lfloor\frac{k\cdot(N-1)}{S}\right\rfloor,k=0,1,\dots,S-1$ (1)
Where $i_{k}$ is the $k$-th index of the sampled data, $N$ is the total size of the dataset, $S$ is the target sample size, and $\left\lfloor x\right\rfloor$ denotes the floor function. The term $\frac{k\cdot(N-1)}{S}$ maps $k$ uniformly to the range $[0,N-1]$, ensuring consistent intervals between sampling points. The floor function ensures valid integer indices.
III-C Knowledge Point and Main Idea Extraction
In this section, the relevant knowledge points are extracted from the sampled dataset. During the extraction process, LLMs autonomously generate knowledge points, which need to meet the following requirements:
  1. 1.
    The generated knowledge points must be relevant to the question.
  2. 2.
    They should focus on summarizing broad concepts or categories of knowledge related to the problem rather than detailed explanations.
  3. 3.
    The generated content must adhere to a predefined format. If it does not, the LLM is re-invoked to generate compliant results.
Following the extraction of knowledge points, the main idea of the question is summarized. The LLM generates the main idea by summarizing the question, adhering to the following two requirements:
  1. 1.
    The main idea must encompass the context of the problem and the core of the answer.
  2. 2.
    The content must be concise and clear, avoiding redundant expressions.
The main idea plays a key role in the subsequent online search and question design stages, ensuring that the newly generated questions align with the core ideas of the static dataset questions.
During the extraction of knowledge points and main ideas, a few-shot strategy is employed, providing examples tailored to different datasets to improve the quality of the LLM-generated content. Overall, the knowledge points depict various aspects of the current question, while the main idea focuses on its core concept. This design ensures consistency in the core ideas while expanding the content of new questions in different directions.
The prompts and few-shot examples for extracting knowledge points and main ideas are detailed in Appendices  A-A,  A-B and  B-A,  B-B.
Below is an example of knowledge point and main idea extraction.
Extract Knowledge Point and Main Idea Example Input Question: Which of the following regular expressions is equivalent to (describes the same set of strings as) (a* + b)*(c + d)? ? A a*(c + d)+ b(c + d). B a*(c + d)* + b(c + d)*. C a*(c + d)+ b*(c + d). D (a + b)*c +(a + b)*d. Answer: D
Example Output Knowledge Point: [”Regular expressions and their operators”, ”Concatenation in regular expressions”, ”Union operator (+) in regular expressions”, ”Kleene star (*) in regular expressions”, ”Equivalence of regular expressions”, ”Pattern matching with regular expressions”] Main Idea: Finding the regular expression equivalent to (a* + b)*(c + d).
III-D Online Search for Knowledge Point Explanations
In this section, we utilize the online search function of large language models to input the problem statement, related knowledge points, and corresponding main ideas of each question in the static dataset. The model then generates detailed explanations for the knowledge points. The generated content needs to meet the following three requirements:
  1. 1.
    Focus on the Knowledge Points: The detailed explanation must be closely centered around the knowledge points provided, and can include background or supplementary information from the question or main ideas, but the focus must remain on the knowledge points.
  2. 2.
    Concise Paragraphs: The output must be a clear and coherent paragraph, avoiding overly long or excessively short content, and must not include paragraph breaks or excessive embellishments.
  3. 3.
    Straightforward: The output must be comprehensive, accurate, and logically clear, focusing on the main topic while avoiding irrelevant information or vague statements.
Each question is associated with a main idea and multiple knowledge points. To increase the flexibility of generation, we designed a knowledge point selection mechanism that controls the number of knowledge points selected. In practical tests, we found that generating a detailed explanation for one knowledge point using the Qwen-Plus API takes about 30 seconds. If explanations are generated for each knowledge point, it would lead to considerable time and cost consumption.
To address this, we introduced a knowledge point selection strategy: we randomly select a specified number of knowledge points from the array of knowledge points for processing. For example, if the parameter is set to 2, and the number of knowledge points exceeds 2, 2 points are selected at random; if there are fewer than 2, all knowledge points are selected. We then traverse the selected knowledge points and generate related explanations based on the problem statement and main ideas.
The online search prompts and few-shot examples are detailed in Appendix  A-C and  B-C. Below is an example of the online search for knowledge point explanations.
Online Search Example Example Output Regular expressions are sequences of characters that define search patterns, primarily used for pattern matching within strings. The fundamental operators in regular expressions include concatenation (placing expressions sequentially), alternation (denoted by ‘—‘ or ‘+‘, representing a choice between expressions), and the Kleene star (denoted by ‘*‘, indicating zero or more repetitions of the preceding element). Understanding these operators is crucial for constructing expressions that accurately represent specific …
III-E Question Design
This section aims to generate two types of new questions by embedding the aforementioned static dataset questions, knowledge points, main ideas, and knowledge point explanations into the prompts: general new questions and Bloom’s Taxonomy-based[41, 42] new questions. The generated questions needs to meet the following requirements:
  1. 1.
    High Relevance: The question, correct option, and answer must be directly based on the provided text and closely related to the main idea.
  2. 2.
    Consistent/Greater/Lesser Difficulty: The complexity of the question should match the reference question’s complexity, be slightly higher, or slightly lower, ensuring that the question requires deep understanding, analysis, and synthesis of details to arrive at the correct answer, rather than being a simple recall of facts.
  3. 3.
    Reasonable and Discriminative Options: The options should be logically clear, avoiding obvious answers. At least one strong distractor should be included to assess the student’s attention to detail and reasoning ability.
  4. 4.
    Avoid External Dependencies: The question and options must be entirely based on the provided text and main idea. If external content is required, it should be embedded directly into the question, avoiding references such as ”According to the text” or similar phrases. Questions should not rely on context that makes them difficult to understand in isolation.
  5. 5.
    Unique Answer: Ensure only one option is correct, with other options being plausible but subtly incorrect or only partially matching the main idea and text content.
  6. 6.
    Assess Deep Understanding: The designed questions can involve inference, comparison, method analysis, application of key concepts, etc., to increase the difficulty.
  7. 7.
    Innovative Question Stem: The question format should not be too similar to the reference questions, demonstrating your ability to innovate, while maintaining rigor and professionalism in the question style.
  8. 8.
    Different Emphasis: The generated questions should focus on different aspects compared to the reference questions. For example, if the reference question focuses on concept definitions, the generated question could focus on application of methods, comparison of key details, analysis of pros and cons, or practical scenario judgment from other angles.
  9. 9.
    Focus on Knowledge Points: The question should be centered around the provided knowledge points, with the core content tightly aligned with the main idea.
The newly generated questions based on Bloom’s taxonomy, in addition to meeting the requirements 1-5 and 9 listed above, must also satisfy the following: Level Coverage: Each question corresponds to a cognitive level, covering all six levels of Bloom’s taxonomy (for specific details, see Appendix  A-E). There should be one question for each level, totaling six questions.
The question design prompts and few-shot examples are detailed in Appendices  A-D,  A-E, and  B-D,  B-E. Below is an example of a new question based on Bloom’s Taxonomy(The complete example can be found in Appendix  C):
Question Design Example Based on Bloom’s Taxonomy Layer: Remembering Question: Which symbol in regular expressions denotes zero or more repetitions of the preceding element? A: ? B: * C: + D: . Answer: B Layer: Creating Question: If you were to create a new regular expression based on the pattern ‘(a* + b)*(c + d)‘ that matches a string starting with zero or more ’a’s, followed by one or more ’b’s, and ending with either ’c’ or ’d’, what would that regular expression be?, A: (a*b+)(c—d) B: (a+b*)(c—d) C: (a* + b+)(c—d) D: (a+b*)(c+d) Answer: C
III-F Question Complexity Control
If no constraints are applied to the LLM’s question generation process, the results tend to produce relatively simple datasets. To solve this issue, we designed a question difficulty adjustment process (as shown in Figure 3) to ensure that the generated dataset’s complexity aligns with that of the static dataset.
Figure 3: Question Difficulty Adjustment Process - Aligning with Static Dataset Example
In the initial phase of generating the dataset, we introduce a complexity alignment mechanism through few-shot prompting to ensure the complexity of the generated data matches expectations. However, in the initial phase, there is a large discrepancy in complexity, so further adjustments are necessary to control the complexity of the dataset within an acceptable error range.
Specifically, we indirectly measure the complexity of the dataset using the accuracy of the preset model (e.g., Qwen-7B-Chat). The newly generated dataset is tested with the preset model, and the accuracy of the model is calculated to derive the deviation in question difficulty:
$\Delta=s_{\text{target}}-s_{\text{gen}}$ (2)
Where $s_{\text{target}}$ represents the desired accuracy (for example, when aligning with the static dataset, it refers to the accuracy of the static dataset on the preset model), and $s_{\text{gen}}$ represents the accuracy of the newly generated dataset on the preset model. We aim for the deviation to satisfy $|\Delta|<5\%$.
Adjustment Strategy:
If $\Delta>5\%$: The generated dataset’s complexity is too high. We then select the questions that the model answered incorrectly from the generated dataset and reconstruct them to simplify the questions.
If $\Delta<-5\%$: The generated dataset’s complexity is too low. We then select the questions that the model answered correctly and reconstruct them to increase the complexity.
After each reconstruction, the newly generated dataset is re-tested, and the above steps are repeated until $|\Delta|<5\%$. Through this process, we can effectively control the complexity of the generated dataset, aligning it with the static dataset or preset target.
The complexity control prompts and few-shot examples are detailed in Appendices  A-F and  B-F.
III-G Quality Control
To ensure that each step of the AdEval process meets quality requirements, we introduced a multi-model voting mechanism for quality assessment. Specifically, multiple LLMs (including Doubao-pro-32k, Qwen-plus, and GPT-4o-mini) are used for quality evaluation. The voting rules are as follows:
  1. 1.
    If at least 2 out of the 3 responses are judged as ”pass,” the question is considered of acceptable quality.
  2. 2.
    If the above condition is not met, the question is deemed to have quality issues and is regenerated.
The quality control prompts are detailed in Appendix  A-G.
IV Experiment
IV-A Experimental Setup
IV-A1 Task and Complexity
The datasets we used include: MMLU [11] and ARC-Challenge [40], with 300 samples generated for each dataset. To reduce the interference of randomness on the evaluation results, each generated dataset was evaluated independently 5 times, while ensuring that the complexity of the generated data aligns with that of the static dataset.
To balance performance and computational costs, we selected Qwen-plus (API) as the primary tool for generating knowledge points, main ideas, and detailed explanations of knowledge points, and used Doubao-pro-32k (API) to align question complexity. The temperature parameter for the large models was set to 0 to avoid interference from randomness during generation.
IV-A2 Dynamic Generation and Few-Shot Settings
During the dynamic question generation process, we adopted a Few-Shot learning strategy. Specifically, knowledge point extraction, main idea extraction and explanation, and question generation all use 2-shot; whereas for Bloom classification question generation, 1-shot was used. We observed that using 2-shot for Bloom classification tasks caused a decrease in generation quality due to the excessive length of the context. Therefore, the 1-shot strategy is more suitable for such tasks.
IV-A3 Model Setup
The experiment selected 9 LLMs from different organizations, each with different sizes and access methods. Specific details can be found in Table I.
IV-A4 Evaluation Metrics
To comprehensively assess the performance of LLMs, this experiment uses the following two core metrics:
1. Accuracy Accuracy is calculated by comparing the model’s output to the standard answers, using the formula:
$P=\frac{T}{T+F}$ (3)
where $P$ is accuracy, $T$ is the number of correct answers, and $F$ is the number of incorrect answers.
2. Perplexity Perplexity measures the model’s ”uncertainty” in generating sequences. The specific calculation process is as follows:
1) Log-Probability of a Single Token
The model outputs the log-probability of each token. The natural logarithm is converted to a base 2 logarithm:
$\log_{2}(P)=\frac{\ln(p)}{\ln(2)}$ (4)
2) Average Log-Probability The average log-probability of all generated tokens is calculated as:
$\text{AvgLogP}=\frac{1}{N}\sum_{i=1}^{N}\log_{2}(p_{i})$ (5)
where $N$ is the number of generated tokens.
3) Perplexity Calculation Based on the average log-probability, perplexity is calculated as:
$\text{PPL}=2^{-\text{AvgLogP}}\log_{2}(p_{i})$ (6)
4) Overall Dataset Perplexity The perplexity of all samples in the dataset is averaged:
$\text{PPL}_{dataset}=\frac{1}{M}\sum_{j=1}^{M}\text{PPL}_{j}$ (7)
where $M$ is the total number of samples in the dataset, and $\text{PPL}_{j}$ is the perplexity of the $j$-th sample.
IV-A5 Fine-Tuning Setup
The fine-tuning process used the LoRA method, with the following specific configuration: the learning rate was set to $5\times 10^{-5}$, the number of training epochs was 2, the batch size was 2, the LoRA rank was 8, the scaling factor was 16, the random dropout rate was 0.1, and the LoRA+ learning rate ratio was 16.
IV-B Data Contamination Fine-Tuning Test
This experiment aims to explore the impact of data contamination on the fine-tuning effect of small models. Two models were compared: one is the clean model (CLN), and the other is the model contaminated by static data (CTM). The experimental results were compared and analyzed using different datasets, quantifying the effects of static data contamination and dynamic data generation methods on model performance.
CLN: Clean model; CTM: Contaminated model with static data; $\Delta$: The accuracy difference between CLN and CTM. Orange indicates $\Delta>10$, magenta indicates $\Delta\in[5,10]$, and blue indicates $\Delta\in[0,5)$; Static: MMLU static data; DataGM (Data Generation Method): Dynamic data generation methods; AdEval: Data dynamically generated by the AdEval method; Auto-Dataset: Data dynamically generated by Auto-Dataset; LatestEval: Data dynamically generated by LatestEval
In Table II, the accuracy from static data on the clean model is used as a baseline, indicated with an underscore (e.g., 60.9±0.8). If the accuracy of a dynamic data generation method for the clean or contaminated model differs from this baseline by less than 5%, it is also indicated with an underscore.
Model Creator Parameters Access
GPT4o OpenAI Undisclosed API
GPT4o OpenAI Undisclosed API
GPT-3.5-turbo OpenAI Undisclosed API
Claude-3-5-haiku Anthropic Undisclosed API
Doubao-pro-32k ByteDance Undisclosed API
glm-4-flash Tsinghua Undisclosed API
Llama-3-3-70B Meta 70B API
Qwen-7B-Chat Alibaba 7B Weights
chatglm3-6b Tsinghua 6B Weights
glm-4-9b-chat Tsinghua 9B Weights
TABLE I: LLMs used in the experiment.
DataGM Qwen2-7B-Chat % Meta-Llama3-8B-Instruct % glm-4-9b-chat %
CLN CTM $\Delta$ CLN CTM $\Delta$ CLN CTM $\Delta$
MMLU
Static 60.9$\pm$0.8 85.2$\pm$1.9 24.3$\uparrow$ 55.5$\pm$1.2 84.4$\pm$3.4 28.9$\uparrow$ 64.7$\pm$1.0 90.3$\pm$1.0 25.6$\uparrow$
AdEval 60.4$\pm$0.6 64.6$\pm$1.6 4.2$\uparrow$ 60.8$\pm$1.0 63.0$\pm$0.7 2.2$\uparrow$ 66.5$\pm$1.0 62.9$\pm$1.1 3.6$\downarrow$
Auto-Dataset 76.5$\pm$0.8 68.4$\pm$0.7 8.1$\downarrow$ 66.0$\pm$1.0 71.0$\pm$1.7 5.0$\uparrow$ 66.0$\pm$2.0 73.2$\pm$0.9 7.2$\uparrow$
LatestEval 84.2$\pm$0.4 86.6$\pm$0.8 2.4$\uparrow$ 79.2$\pm$9 79.6$\pm$0.4 0.2$\uparrow$ 77.5$\pm$0.3 81.3$\pm$0.5 3.8$\uparrow$
ARC
Static 83.0$\pm$1.2 99.0$\pm$0.3 16.0$\uparrow$ 73.1$\pm$0.9 96.3$\pm$0.4 23.2$\uparrow$ 83.8$\pm$1.3 94.6$\pm$0.4 10.8$\uparrow$
AdEval 81.4$\pm$0.8 81.1$\pm$1.3 0.3$\downarrow$ 74.6$\pm$0.5 73.8$\pm$0.7 0.8$\downarrow$ 82.7$\pm$1.3 82.7$\pm$0.3 0.0
Auto-Dataset 86.4$\pm$0.7 90.5$\pm$0.8 4.1$\uparrow$ 82.0$\pm$0.0 88.8$\pm$0.5 6.8$\uparrow$ 89.7$\pm$1.0 89.5$\pm$1.5 0.2$\downarrow$
LatestEval 88.4$\pm$0.4 92.6$\pm$0.2 4.2$\uparrow$ 86.0$\pm$0.4 88.7$\pm$0.7 1.7$\uparrow$ 90.7$\pm$0.4 87.8$\pm$0.3 2.9$\downarrow$
TABLE II: Impact of Data Contamination on Accuracy across Different Models and Data Generation Methods
The following conclusions can be drawn from Table II:
  1. 1.
    Significant Improvement in Model Performance with Static Data Contamination: Fine-tuning with static data contamination (Static) significantly improves the accuracy of all models, with $\Delta$ values exceeding 10%, most exceeding 20%. This shows that static data contamination has a strong positive impact on model performance. For example, in the MMLU dataset, the $\Delta$ values for static data contamination are 24.3%, 28.9%, and 25.6%, significantly higher than other dynamic data generation methods.
  2. 2.
    Minimal Impact of Contaminated Data on Dynamic Data: Among the three dynamic data generation methods, AdEval and LatestEval maintain $\Delta$ values within 5%, indicating that dynamic data is more stable under contamination. For example, in the MMLU dataset, AdEval’s $\Delta$ value is 4.2%, LatestEval’s $\Delta$ value is 2.4%; in the ARC dataset, AdEval’s $\Delta$ value is 0.3%, LatestEval’s $\Delta$ value is 4.2%. In contrast, Auto-Dataset shows higher $\Delta$ values in some experiments, such as 8.1% in the MMLU dataset, indicating larger fluctuations.
  3. 3.
    AdEval’s Data Difficulty Consistent with Static Data: The data generated by AdEval is consistent with static data in terms of difficulty. Whether the model is contaminated or not, the accuracy of the data generated by AdEval differs by less than 5% from the baseline accuracy of static data on the clean model. For example, in the MMLU and ARC-Challenge datasets, the accuracy of the data generated by AdEval closely matches static data, with the difficulty significantly higher than that of LatestEval and Auto-Dataset. In contrast, the data generated by LatestEval and Auto-Dataset tends to be simpler in some experiments.
  4. 4.
    Limitations of Auto-Dataset and LatestEval: The data generated by Auto-Dataset and LatestEval is relatively simpler in some experiments. For example, in the MMLU dataset with the clean model, the accuracy of static data is 60.9%, while the data generated by Auto-Dataset and LatestEval show accuracy rates of 76.5% and 84.2%, respectively.
In conclusion, regardless of model contamination, AdEval can effectively evaluate model capabilities, and the data generated by AdEval is highly consistent in difficulty with static datasets.
IV-C Data Reconstruction
This experiment demonstrates how AdEval controls data complexity through data reconstruction. We selected the data reconstruction process of AdEval on the MMLU dataset, using the Qwen2-7B-Chat model. AdEval progressively adjusts the complexity of the generated dataset through multiple rounds of data reconstruction to make it largely consistent with the complexity of static data. Figure 4 shows the accuracy changes of the clean model (CLN) and contaminated model (CTM), with the baseline accuracy of static data on the clean model shown by the red horizontal dashed line for comparison.
Figure 4: Question Difficulty Adjustment Process - Aligning with Static Dataset Example
As shown in Figure 4, with each round of data reconstruction, the complexity of the data generated by AdEval gradually increases. This is reflected by the changes in model accuracy on the dataset (the lower the accuracy, the higher the complexity). The experimental results show that the accuracy of the data generated by AdEval gradually decreases from its initial higher value, indicating that AdEval effectively increases the complexity of the generated data through data reconstruction. After a certain number of rounds of reconstruction, both the clean model (CLN) and the contaminated model (CTM) show an accuracy that is nearly identical to the baseline accuracy of the static data (indicated by the red horizontal dashed line).
In conclusion, whether the model is contaminated or not, AdEval can effectively regulate the complexity of the data through multiple rounds of reconstruction, ultimately generating a dataset with a complexity level consistent with that of the static data.
IV-D Bloom’s Cognitive Hierarchy
Figure 5: Question Difficulty Adjustment Process - Aligning with Static Dataset Example
MMLU % ARC-Challenge %
CLN CTM $\Delta$ CLN CTM $\Delta$
Static 1.31 1.14 0.17$\downarrow$ 1.24 1.01 0.23$\downarrow$
AdEaval 1.32 1.23 0.09$\downarrow$ 1.30 1.29 0.01$\downarrow$
Auto-dataset 1.22 1.16 0.06$\downarrow$ 1.21 1.03 0.18$\downarrow$
Latesteval 1.20 1.07 0.13$\downarrow$ 1.20 1.06 0.14$\downarrow$
TABLE III: Impact of Data Contamination on Accuracy across Different Models and Data Generation Methods
The radar chart (Figure 5) generated by AdEval reveals that the difficulty of dynamically generated questions in the MMLU dataset is significantly higher than that in the ARC-Challenge dataset. Despite the overall difficulty differences between the two datasets, the internal difficulty distribution within each dataset remains consistent. For example, questions in the ”Creating” and ”Applying” levels are more challenging in both datasets, while questions in the ”Understanding” and ”Evaluating” levels are less difficult. This demonstrates AdEval’s stability in generating questions across different cognitive levels.
Additionally, when comparing the results of Qwen-7B-Chat-CLN (clean model) and Qwen-7B-Chat-CTM (contaminated model), although data contamination slightly affects certain cognitive levels, the overall impact on the model remains minor, and the difficulty stability is maintained.
IV-E Perplexity
This experiment evaluated the perplexity of the Qwen-7B-Chat clean model (CLN) and contaminated model (CTM) across multiple datasets. The results are shown in Table 3, with $\Delta$ values highlighted in different colors: orange for $\Delta>0.2$, magenta for $\Delta\in[0.1,0.2]$, and blue for $\Delta\in[0,0.1]$.
From Table III, it can be observed that static data fine-tuning (data contamination) significantly reduces model perplexity, with $\Delta$ values around 0.2. For example, in the MMLU dataset, the $\Delta$ for static data fine-tuning is 0.17, while in the ARC-Challenge dataset, it is 0.23.
Among dynamic data generation methods, AdEval exhibits $\Delta$ values below 0.1, indicating minimal and stable reductions. In contrast, LatestEval shows $\Delta$ values between 0.1 and 0.2, while Auto-Dataset displays a wider range, with $\Delta$ spanning from 0.05 to 0.2, showing more significant fluctuations.
It is noteworthy that both static and dynamic data fine-tuning reduce the model’s perplexity. This is because all generated data belong to the same question types as static data. Although the specific questions differ, fine-tuning familiarizes the model with these question types, leading to an overall decrease in perplexity. Among all methods, AdEval demonstrates the smallest and most stable reduction in perplexity.
IV-F Performance Evaluation
We evaluated the LLMs listed in Table 1 using AdEval.
This experiment tested various large models on dynamically generated data using different methods. A represents testing using only dynamic data questions, while B indicates adding the corresponding static data questions into the prompts of A. $\Delta$ represents the difference between A and B, with the results color-coded as follows: orange for $\Delta>5$, magenta for $\Delta\in[2,5]$, and blue for $\Delta\in[0,2)$.
Note: Due to stricter safety settings, the glm-4-flash model refused to answer approximately 5% of questions, causing some fluctuation in its results.
AdEval % Auto-Dataset % LatestEval %
A B $\Delta$ A B $\Delta$ A B $\Delta$
gpt-3.5-turbo 67.0$\pm$0.8 65.9$\pm$1.9 1.1$\downarrow$ 65.8$\pm$2.5 70.4$\pm$1.4 4.6$\uparrow$ 80.6$\pm$0.8 84.7$\pm$0.0 4.1$\uparrow$
gpt-4o-mini 74.8$\pm$0.0 74.4$\pm$0.6 0.4$\downarrow$ 81.0$\pm$0.7 82.9$\pm$0.9 1.9$\uparrow$ 87.3$\pm$0.5 91.4$\pm$0.4 4.1$\uparrow$
Doubao-pro-32k 84.0$\pm$0.3 84.6$\pm$0.3 0.2$\uparrow$ 85.0$\pm$0.7 87.4$\pm$0.4 2.4$\uparrow$ 88.9$\pm$0.0 94.6$\pm$0.1 5.7$\uparrow$
qwen-plus 79.4$\pm$0.2 79.9$\pm$0.7 0.5$\uparrow$ 84.4$\pm$0.4 90.0$\pm$0.7 5.6$\uparrow$ 90.0$\pm$0.2 96.1$\pm$0.3 9.6$\uparrow$
claude-3-5-haiku 64.8$\pm$0.6 67.7$\pm$1.0 2.9$\uparrow$ 75.2$\pm$1.8 76.0$\pm$2.0 0.8$\uparrow$ 84.1$\pm$1.6 88.0$\pm$0.5 3.9$\uparrow$
glm-4-flash 66.1$\pm$0.3 71.4$\pm$0.4 5.3$\uparrow$ 69.0$\pm$1.3 72.8$\pm$1.8 3.8$\uparrow$ 78.5$\pm$0.8 86.5$\pm$0.1 8.0$\uparrow$
TABLE IV: Performance Comparison of LLMs Using Dynamic Data
From Table IV, it can be observed that:
  1. 1.
    The impact of static data on dynamic datasets generated by the three methods: AdEval ¡ Auto-Dataset ¡ LatestEval.
  2. 2.
    Difficulty of dynamic datasets generated by the three methods: AdEval ¿ Auto-Dataset ¿ LatestEval.
IV-G Quality Control
We conducted quality checks on questions and answers generated by AdEval, Auto-Dataset, and LatestEval. Three large language models (Doubao-pro-32k, Qwen-plus, and GPT-4o-mini) voted to assess correctness. The rule was that if at least two of the three models judged an answer to be correct, the question was marked as correct; otherwise, it was marked as incorrect.
The table below shows the error rates for dynamic datasets generated by the three methods: AdEval Auto-Dataset LatestEval MMLU ARC-Challenge
AdEval Auto-Dataset LatestEval
MMLU 5.8% 9.0% 7.5%
ARC-Challenge
TABLE V: Error Rates of Dynamic Data Generation Methods
From Table V, it can be observed that: AdEval has the lowest error rate, indicating superior accuracy in question generation compared to Auto-Dataset and LatestEval.
V Conclusion and Discussion
The proposed AdEval method achieves remarkable results in addressing the data contamination problem in large language model (LLM) evaluation. By dynamically generating evaluation data consistent with static data, AdEval effectively mitigates the impact of data contamination while precisely controlling question complexity, ensuring the generated datasets share the same distribution as static datasets. Experimental results demonstrate that AdEval outperforms other dynamic data generation methods on the MMLU and ARC-Challenge datasets, particularly in controlling data complexity and maintaining quality stability.
Under the influence of data contamination, static data fine-tuning effectively enhances model performance, while dynamic data generation methods exhibit varying degrees of resilience to contamination. Compared to Auto-Dataset and LatestEval, AdEval better controls the complexity and quality of generated questions while minimizing the impact of data contamination on the model. Particularly in complexity control, AdEval shows significant potential in applications across different cognitive levels and data domains, providing a more stable and reliable solution for LLM evaluation.
Future research could explore the scalability and diversity of AdEval, especially by validating and applying it to more types of datasets. Key directions include enhancing the diversity and adaptability of generated data, optimizing generation algorithms, and reducing potential biases and errors during the generation process.
Appendix A Prompts
A-A Knowledge Point Extraction - Prompt
You are a professional expert in summarizing knowledge points from questions. Please summarize the relevant knowledge points from the given multiple-choice question.
Requirements:
1. Provide professional knowledge points related to the question.
2. Focus on summarizing broad concepts or categories of knowledge related to the question, rather than specific detailed explanations.
3. Strictly follow the format below and do not output any other content:
[“knowledge1”, “knowledge2” , “knowledge3”, …]
{{few-shot}}
Please summarize the following multiple-choice question:
{{choiceQ}}
The term {{few-shot}} refers to few-shot learning, with specific details provided in Appendix  B-A. {{choiceQ}} refers to the questions in the dataset. Both of these are embedded in the prompt and provided as input to the large language model.
A-B Main Idea Extraction - Prompt
You are a professional expert in summarizing the main idea of questions. Please summarize the main idea based on the input question content. Requirements:
1. Provide the main idea related to the question, covering the background of the question and the core of the answer.
2. Ensure the main idea is concise and clear, avoiding redundant expressions.
3. Output only the main idea without any additional content.
{{few-shot}}
Please summarize the main idea of the following question:
{{choiceQ}}
A-C Online Search - Prompt
You are an expert in explaining knowledge points. Based on the provided question, knowledge point, and main idea, use online resources to search for and provide a detailed explanation closely related to the core of the knowledge point. Please strictly follow these requirements for your output:
1. Focus on the Knowledge Point: The explanation must revolve tightly around the input knowledge point. While the question and main idea can provide background or supplementary context, the knowledge point must remain the core focus.
2. Concise Paragraph: Present the explanation in a single, clear, and coherent paragraph. Avoid excessive length or overly brief statements, and do not include additional structuring or embellishments. 3. Direct and Comprehensive: Ensure the explanation is accurate, logically sound, and centered on the topic, avoiding irrelevant information or ambiguous statements.
{{few-shot}}
Please search online and provide a single-paragraph detailed explanation of the knowledge point, referencing the question and main idea where relevant.
Actual Input Question:
{{choiceQ}}
Actual Input Knowledge Point:
{{kn}}
Actual Input Main Idea:
{{purport}}
{{kn}} and {{purport}} represent the knowledge points and main ideas extracted in  A-A and  A-B.
A-D Question Design - Prompt
You are a professional question designer specializing in creating challenging multiple-choice questions. Based on the provided reference question, knowledge points, main ideas, and text content, please generate one multiple-choice question and its answer. Ensure the question’s difficulty is not lower than the reference question, and preferably slightly higher. The question should test students’ in-depth understanding, detailed analysis, and comprehensive judgment skills. The multiple-choice question must meet the following requirements:
1. High Relevance: The question, correct option, and answer must be directly based on the provided text and closely related to the main idea.
2. Consistent or Higher Difficulty: The complexity of the question must match or exceed that of the reference question. Avoid simple factual recitations; instead, require students to achieve a deep understanding and analyze details to deduce the correct answer.
3. Reasonable and Distinguishable Options: The options must be logically clear, avoiding overly obvious correct answers. Include at least one strong distractor to test detailed reading and reasoning skills.
4. Avoid external dependency: The question and its options must be entirely based on the provided text content and main idea. If quoting content is necessary, directly embed the relevant text in the question rather than using references like ”according to the passage.” Avoid creating questions that depend on context not provided, such as: ”Which of the following best describes the purpose of penetration testing as described in the text?”
5. Unique Answer: Ensure only one option is correct, while other options remain plausible but subtly incorrect or mismatched with the main idea and text content.
6. Deep Understanding: The question can involve inference, comparison, method analysis, application of key concepts, etc., to increase the difficulty level.
7. Innovative Question Stem: The question format should differ significantly from the reference question, showcasing creativity while maintaining a rigorous and professional style.
8. Different Focus: The question’s focus should differ from the reference question. For example, if the reference question emphasizes concept definition, the generated question could focus on method application, key detail comparison, pros and cons analysis, or judgment in practical scenarios.
9. Focus on knowledge points: The question must center around the provided knowledge points, with its core content closely aligned with the main idea.
10. Strict Output Format: The output must strictly follow the JSON structure below, without any additional content:
    "Question": "Question content",
    "A": "Option content",
    "B": "Option content",
    "C": "Option content",
    "D": "Option content",
    "Answer": "A/B/C/D"
{{few-shot}}
Below are the provided reference question, knowledge points, main idea, and text content. Please generate a multiple-choice question and answer that meets the above requirements. Remember, do not make the question similar to the reference question, but ensure the difficulty level is consistent with or even higher than the reference question.
Reference question:
{{choiceQ}}
Knowledge Point:
{{kn}}
Main Idea:
{{purport}}
Text Content:
{{KNexplain}}
{{KNexplain}} refers to the detailed explanations of the knowledge points extracted in  A-C.
A-E Bloom’s Taxonomy Question Design - Prompt
You are a professional question designer specializing in creating challenging multiple-choice questions. Based on the provided reference question, knowledge points, main ideas, and text content, please generate six multiple-choice questions and answers corresponding to Bloom’s six levels of cognitive learning. Ensure the question’s difficulty is not lower than the reference question, and preferably slightly higher. The question should test students’ in-depth understanding, detailed analysis, and comprehensive judgment skills.
Let me first explain what Bloom’s six cognitive levels are:
Introduction to Bloom’s Six Cognitive Levels:
Level 1: Remembering
This involves recognizing and recalling concepts and knowledge, storing them in the brain, and retrieving them when needed. Examples include memorizing vocabulary, poetry, or definitions. This level involves identifying concrete or abstract knowledge and can pertain to factual, conceptual, procedural, or metacognitive knowledge. While mechanical, it is a foundational step for learning and solving more complex problems.
Level 2: Understanding
Understanding means grasping the essence of information or knowledge, though not deeply or thoroughly—just a preliminary understanding. When learners connect ”new” knowledge to existing knowledge, they achieve understanding. Examples include the classic Feynman technique, where teaching others reinforces personal understanding. Understanding occurs when new information integrates with existing cognitive frameworks.
Understanding includes translating, interpreting, and inferring:
Interpreting involves explaining or summarizing information in one’s own words. Translating involves expressing learned content in different forms, such as using diagrams to explain a concept.
Inferring involves predicting trends or developments based on learned knowledge.
Level 3: Applying
This level focuses on applying learned concepts, principles, or rules in new situations or solving real-world problems. It involves correctly using abstract ideas in appropriate contexts without explicit problem-solving instructions. Applications rely on remembering and understanding as foundational, including applying concepts, principles, methods, and theories, such as using the Pythagorean theorem to solve geometry problems.
Level 4: Analyzing
Analyzing involves breaking down complex knowledge into its components and understanding the relationships between them. It requires deconstructing material to identify essential elements and their relationships, making the material’s organization and structure clearer. Tasks include identifying, analyzing, and recognizing the interrelationships and structure among concepts.
For instance, discovering how Factor A changes alongside Factor B by analyzing correlations, causal relationships, mediating variables, and moderators.
Level 5: Evaluating
This level involves making judgments about the value of something based on rational and persuasive reasoning rather than intuition or observation. It requires synthesizing internal and external evidence and making objective, well-supported assessments. For example, legal debates where both sides cite laws to support their arguments exemplify evaluation.
Level 6: Creating
This involves reorganizing learned knowledge or generating new information to form a new whole, such as proposing new ideas, solutions, or designs. It requires analyzing and synthesizing components to form a new structure. Creativity emphasizes innovation and problem-solving, breaking conventional thought patterns and achieving the ultimate learning goal.
Creating is the highest level as it challenges learners to develop new knowledge structures, highlighting innovation and originality.
Multiple-Choice Question Requirements:
1. High Relevance: The question, correct option, and answer must be directly based on the provided text and closely related to the main idea.
2. Consistent or Higher Difficulty: The complexity of the question must match or exceed that of the reference question. Avoid simple factual recitations; instead, require students to achieve a deep understanding and analyze details to deduce the correct answer.
3. Reasonable and Distinguishable Options: The options must be logically clear, avoiding overly obvious correct answers. Include at least one strong distractor to test detailed reading and reasoning skills.
4. Avoid external dependency: The question and its options must be entirely based on the provided text content and main idea. If quoting content is necessary, directly embed the relevant text in the question rather than using references like ”according to the passage.” Avoid creating questions that depend on context not provided, such as: ”Which of the following best describes the purpose of penetration testing as described in the text?”
5. Unique Answer: Ensure only one option is correct, while other options remain plausible but subtly incorrect or mismatched with the main idea and text content.
6. Focus on knowledge points: The question must center around the provided knowledge points, with its core content closely aligned with the main idea.
7. Coverage of Levels: Each question must correspond to a single cognitive level, covering all six levels of Bloom’s taxonomy, with one question per level. A total of six questions is required.
8. Strict Output Format: The output must strictly follow the JSON structure below, without any additional content:
    "Layer": "Cognitive Level",
    "Question": "Question content",
    "A": "Option content",
    "B": "Option content",
    "C": "Option content",
    "D": "Option content",
    "Answer": "A/B/C/D"
{{few-shot}}
Below are the provided reference question, knowledge points, main idea, and text content. Please generate Bloom’s six cognitive-level multiple-choice questions that meets the above requirements. Remember, do not make the question similar to the reference question, but ensure the difficulty level is consistent with or even higher than the reference question.
Reference question:
{{choiceQ}}
Knowledge Point:
{{kn}}
Main Idea:
{{purport}}
Text Content:
{{KNexplain}}
A-F Complexity Control - Prompt
You are a professional question reconstruction expert, specializing in enhancing question complexity and designing challenging multiple-choice questions.
Please redesign a multiple-choice question based on the provided original question, current question, knowledge points, main idea, and text content. The goal is to improve students’ in-depth understanding, detailed analysis, and comprehensive judgment skills.
Multiple-Choice Question Design Requirements:
1. High Relevance: The stem, correct option, and distractors must be strictly based on the provided text content. Ensure the question is closely related to the main idea, and the answer can be explicitly supported by the text.
2. Increased Difficulty: The question should be more complex than the current one. Avoid simple factual repetition; instead, require students to deeply understand, analyze details, and make connections to arrive at the correct answer.
3. Reasonable and Distinguishable Options: The options must be logically clear, avoiding overly obvious correct answers. Include at least one strong distractor to test detailed reading and reasoning skills.
4. Avoid external dependency: The question and its options must be entirely based on the provided text content and main idea. If quoting content is necessary, directly embed the relevant text in the question rather than using references like ”according to the passage.” Avoid creating questions that depend on context not provided, such as: ”Which of the following best describes the purpose of penetration testing as described in the text?”
5. Unique Answer: Ensure only one option is correct, while other options remain plausible but subtly incorrect or mismatched with the main idea and text content.
6. Deep Understanding: The question can involve inference, comparison, method analysis, application of key concepts, etc., to increase the difficulty level.
7. Innovative Question Stem: The question format should differ significantly from the original question, showcasing creativity while maintaining a rigorous and professional style.
8. Different Focus: The question’s focus should differ from the original question. For example, if the original question emphasizes concept definition, the generated question could focus on method application, key detail comparison, pros and cons analysis, or judgment in practical scenarios.
9. Focus on knowledge points: The question must center around the provided knowledge points, with its core content closely aligned with the main idea. 10. Strict Output Format: The output must strictly follow the JSON structure below, without any additional content:
    "Question": "Question content",
    "A": "Option content",
    "B": "Option content",
    "C": "Option content",
    "D": "Option content",
    "Answer": "A/B/C/D"
{{few-shot}}
The following includes the original question, current question, knowledge points, main idea, and text content. Please generate a multiple-choice question that meets the above requirements and provide an answer. Be sure not to make it similar to the original question, and ensure that the difficulty of the newly generated question is higher than that of the current question.
Original question:
{{choiceQ}}
current question:
{{choiceQCurrent}}
Knowledge Point:
{{kn}}
Main Idea:
{{purport}}
Text Content:
{{KNexplain}}
{{choiceQCurrent}} refers to the questions dynamically generated in  A-D and  A-E.
A-G Quality Control - Prompt
A-G1 Question Answer Checking Prompt
You are a professional multiple-choice question proofreading expert, specializing in verifying the accuracy of questions and their answers.
I will provide a multiple-choice question and its answer. Please strictly follow the rules below to make your judgment and respond only in numeric form:
1. If the answer is incorrect, respond with ”0”.
2. If the answer is correct, respond with ”1”.
3. Your response must be either 0/1, with no additional content.
Example Input:
[…]
Example Output:
0
Below is a multiple-choice question and its answer. Please judge according to the rules and respond only with 0/1:
{{choiceQ}}
Appendix B Few-Shot
B-A Knowledge Point Extraction - Few-Shot
Example Input1:
[…]
Example Output1:
[”Regular expressions and their operators”, ”Concatenation in regular expressions”, ”Union operator (+) in regular expressions”, ”Kleene star (*) in regular expressions”, ”Equivalence of regular expressions”, ”Pattern matching with regular expressions”]
Example Input2:
”Question: A certain pipelined RISC machine has 8 general-purpose registers R0, R1, . . . , R7 and supports the following operations. ADD Rs1, Rs2, Rd Add Rs1 to Rs2 and put the sum in Rd MUL Rs1, Rs2, Rd Multiply Rs1 by Rs2 and put the product in Rd
[…]
Example Output2:
[”RISC architecture and instruction pipelining”, ”Data hazards and their impact on instruction execution”, ”Register allocation and usage in assembly language”, ”Instruction scheduling to minimize clock cycles”, ”Arithmetic operations in assembly language”]
B-B Main Idea Extraction - Few-Shot
Example Input1:
[…]
Example Output1:
Finding the regular expression equivalent to (a* + b)*(c + d).
Example Input2:
[…]
Example Output2:
Determining the minimum number of clock cycles required to compute the expression $AB+ABC+BC$ on a pipelined RISC machine, considering operation latencies, data dependencies, and constraints on register modification.
B-C Online Search - Few-Shot
Example Input1:
Question:
[…]
Knowledge Point:
Regular expressions and their operators
Main Idea: Finding the regular expression equivalent to (a* + b)*(c + d).
Example Output1:
Regular expressions are sequences of characters that define search patterns, primarily used for pattern matching within strings. The fundamental operators in regular expressions include concatenation (placing expressions sequentially), alternation (denoted by ‘—‘ or ‘+‘, representing a choice between expressions), and the Kleene star (denoted by ‘*‘, indicating zero or more repetitions of the preceding element). Understanding these operators is crucial for constructing expressions that accurately represent specific string patterns. For instance, the expression ‘(a—b)*‘ denotes any combination of the characters ’a’ and ’b’, including the empty string, while ‘a*‘ matches any number of consecutive ’a’ characters, including none. Mastery of these operators enables the creation of complex patterns for efficient text processing and matching tasks.
Example Input2:
[…]
Knowledge Point:
Register allocation and usage in assembly language
Main Idea:
Determining the minimum number of clock cycles required to compute the expression $AB+ABC+BC$ on a pipelined RISC machine, considering operation latencies, data dependencies, and constraints on register modification.
Example Output2:
In assembly language programming, efficient register allocation and usage are crucial for optimizing performance, especially in pipelined RISC architectures. Registers are limited, high-speed storage locations within the CPU that hold data for immediate processing, reducing the need to access slower main memory. Effective register allocation involves assigning these registers to frequently used variables and intermediate results, minimizing memory access and enhancing execution speed. This process can be complex due to the limited number of registers and the need to manage data dependencies and instruction latencies. Techniques such as graph coloring are often employed to allocate registers efficiently by modeling interference between variables and ensuring that no two simultaneously live variables share the same register. Additionally, understanding the specific roles and constraints of registers in a given architecture is essential, as certain instructions may require operands to reside in specific registers. By carefully managing register allocation and usage, assembly language programmers can significantly improve the efficiency and performance of their code.
B-D Question Design - Few-Shot
[…]
Knowledge Point: Regular expressions and their operators
Main Idea: Finding the regular expression equivalent to (a* + b)*(c + d).
Text Content: Regular expressions are sequences of characters that define search patterns, primarily used for pattern matching within strings. The fundamental operators in regular expressions include concatenation (placing expressions sequentially), alternation (denoted by ‘—‘ or ‘+‘, representing a choice between expressions), and the Kleene star (denoted by ‘*‘, indicating zero or more repetitions of the preceding element). Understanding these operators is crucial for constructing expressions that accurately represent specific string patterns. For instance, the expression ‘(a—b)*‘ denotes any combination of the characters ’a’ and ’b’, including the empty string, while ‘a*‘ matches any number of consecutive ’a’ characters, including none. Mastery of these operators enables the creation of complex patterns for efficient text processing and matching tasks.
Example Output1:
    "Question": "Which of the following
    regular expressions correctly represents
    all strings that start with zero or more
    repetitions of either ’x’ or ’y’, followed
    by exactly two repetitions of either ’m’ or
    ’n’, and end with a single occurrence of
    either ’p’ or ’q’?",
    "A": "(x|y)*((m|n)(m|n))(p|q)",
    "B": "((x|y)*(m|n))^2(p|q)",
    "C": "(x* + y*)((m + n)(m + n))(p + q)",
    "D": "(x|y)^2((m|n)(m|n))(p|q)",
    "Answer": "A"
[…]
Knowledge Point: Register allocation and usage in assembly language
Main Idea: Determining the minimum number of clock cycles required to compute the expression $AB+ABC+BC$ on a pipelined RISC machine, considering operation latencies, data dependencies, and constraints on register modification.
Text Content: In assembly language programming, efficient register allocation and usage are crucial for optimizing performance, especially in pipelined RISC architectures. Registers are limited, high-speed storage locations within the CPU that hold data for immediate processing, reducing the need to access slower main memory. Effective register allocation involves assigning these registers to frequently used variables and intermediate results, minimizing memory access and enhancing execution speed. This process can be complex due to the limited number of registers and the need to manage data dependencies and instruction latencies. Techniques such as graph coloring are often employed to allocate registers efficiently by modeling interference between variables and ensuring that no two simultaneously live variables share the same register. Additionally, understanding the specific roles and constraints of registers in a given architecture is essential, as certain instructions may require operands to reside in specific registers. By carefully managing register allocation and usage, assembly language programmers can significantly improve the efficiency and performance of their code.
Example Output2:
    "Question": "Given the same pipelined
    RISC machine and the expression \( AB +
    ABC + BC \), where A, B, and C are in
    registers R0, R1, and R2 respectively,
    and assuming no register modification
    is allowed, which sequence of operations
    would result in the minimum number of
    clock cycles, and what is that minimum
    number?",
    "A": "Perform MUL R0, R1, R3; MUL R1, R2,
    R4; ADD R3, R4, R5; ADD R5, R2, R6; 7
    cycles",
    "B": "Perform MUL R0, R1, R3; ADD R3, R1,
    R4; MUL R1, R2, R5; ADD R4, R5, R6; 7
    cycles",
    "C": "Perform MUL R0, R1, R3; MUL R3, R2,
    R4; ADD R3, R2, R5; ADD R4, R5, R6; 6
    cycles",
    "D": "Perform MUL R0, R1, R3; MUL R1, R2,
    R4; ADD R3, R4, R5; ADD R5, R1, R6; 6
    cycles",
    "Answer": "C"
Finally, remember not to output in the following format (do not begin with “‘json and end with “‘).
‘‘‘json
    "Question": "Which of the following
    regular expressions correctly represents
    all strings that can be formed by zero or
    more repetitions of the strings made up of
    ’a’ or ’b’, followed by exactly one
    occurrence of either ’c’ or ’d’?",
    "A": "(a|b)*(c|d)",
    "B": "(a*b*+b*a*)(c|d)",
    "C": "(a+b)*(c|d)+",
    "D": "((a+b)*c) + ((a+b)*d)",
    "Answer": "A"
B-E Bloom’s Taxonomy Question Design - Few-Shot
Example Input1:
Reference question:
    When trying to construct a regular
expression to match strings that start
with an optional single occurrence of
either ’e’ or ’f’, followed by zero or
more repetitions of a group containing
either ’g’ or ’h’ and then either ’i’
or ’j’, and end with exactly three
repetitions of either ’k’ or ’l’, which
of the following options would be the
correct regular expression?
    A (e|f)?((g|h)*(i|j))(k|l)^3
    B ((e|f)+(g|h))*(i|j)(k|l)^3
    C (e|f)(g|h)*(i|j)^3(k|l)^3
    D (e|f)*((g|h)+(i|j))(k|l)^3
    Answer:A
Knowledge Point:
Regular expressions and their operators
Main Idea:
Finding the regular expression equivalent to (a* + b)*(c + d).
Text Content:
Regular expressions are sequences of characters that define search patterns, primarily used for pattern matching within strings. The fundamental operators in regular expressions include concatenation (placing expressions sequentially), alternation (denoted by ‘—‘ or ‘+‘, representing a choice between expressions), and the Kleene star (denoted by ‘*‘, indicating zero or more repetitions of the preceding element). Understanding these operators is crucial for constructing expressions that accurately represent specific string patterns. For instance, the expression ‘(a—b)*‘ denotes any combination of the characters ’a’ and ’b’, including the empty string, while ‘a*‘ matches any number of consecutive ’a’ characters, including none. Mastery of these operators enables the creation of complex patterns for efficient text processing and matching tasks.
Example Output1:
    "Layer": "Remembering",
    "Question": "Which symbol in regular
    expressions denotes zero or more
    repetitions of the preceding element?",
    "A": "?",
    "B": "*",
    "C": "+",
    "D": ".",
    "Answer": "B"
    "Layer": "Understanding",
    "Question": "In the context of regular
    expressions, what does the expression
    ‘(a|b)*‘ represent?",
    "A": "Any string that contains only ’a’
    repeated any number of times",
    "B": "Any string that contains only ’b’
    repeated any number of times",
    "C": "Any combination of ’a’ and ’b’,
    including the empty string",
    "D": "Any string that starts with ’a’
    or ’b’ and ends with ’a’ or ’b’",
    "Answer": "C"
    "Layer": "Applying",
    "Question": "Given the regular
    expression ‘(a* + b)*(c + d)‘, which of
    the following regular expressions
    correctly applies the same pattern?",
    "A": "(a|b)*(c|d)",
    "B": "(a+b)*(c+d)",
    "C": "(a*b)*(c+d)",
    "D": "(a+b)*(c|d)",
    "Answer": "D"
    "Layer": "Analyzing",
    "Question": "How does the regular
    expression ‘(a|b)*‘ differ from ‘a*‘
    in terms of the possible strings it
    can match?",
    "A": "They match exactly the same set of
    strings",
    "B": "‘(a|b)*‘ can match strings with ’b’,
    while ‘a*‘ cannot",
    "C": "‘a*‘ can match strings with ’b’,
    while ‘(a|b)*‘ cannot",
    "D": "They both match strings with any
    number of ’a’s, including ’b’s",
    "Answer": "B"
    "Layer": "Evaluating",
    "Question": "Considering the regular
    expression ‘(a* + b)*(c + d)‘, evaluate
    which of the following statements best
    justifies the use of the alternation
    operator ‘|‘ over the concatenation
    operator ‘+‘ in this context?",
    "A": "The ‘|‘ operator allows for a
    choice between ’a’ and ’b’, which is
    necessary for matching a pattern that
    could start with either ’a’ or ’b’",
    "B": "The ‘+‘ operator is used for one or
    more repetitions, which is not suitable
    for optional elements",
    "C": "The ‘|‘ operator is more efficient
    for matching a single character",
    "D": "The ‘|‘ operator is used to match
    either ’a’ or ’b’, but it is not
    necessary for the pattern to be valid",
    "Answer": "A"
    "Layer": "Creating",
    "Question": "If you were to create a new
    regular expression based on the pattern
    ‘(a* + b)*(c + d)‘ that matches a string
    starting with zero or more ’a’s, followed
    by one or more ’b’s, and ending with either
    ’c’ or ’d’, what would that regular
    expression be?",
    "A": "(a*b+)(c|d)",
    "B": "(a+b*)(c|d)",
    "C": "(a* + b+)(c|d)",
    "D": "(a+b*)(c+d)",
    "Answer": "C"
Finally, remember not to output in the following format (do not begin with “‘json and end with “‘).
‘‘‘json
    "Layer": "Remembering",
    "Question": "Which symbol in regular
    expressions denotes zero or more
    repetitions of the preceding element?",
    "A": "?",
    "B": "*",
    "C": "+",
    "D": ".",
    "Answer": "B"
B-F Complexity Control - Few-Shot
Example Input1:
Original question:
[…]
current question:
Which of the following regular expressions correctly represents all strings that start with zero or more repetitions of either ’x’ or ’y’, followed by exactly two repetitions of either ’m’ or ’n’, and end with a single occurrence of either ’p’ or ’q’?
A (x|y)*((m|n)(m|n))(p|q) B ((x|y)*(m|n))
^2(p|q) C (x* + y*)((m + n)(m + n))(p +
[…]
Knowledge Point:
Regular expressions and their operators
Main Idea:
Finding the regular expression equivalent to (a* + b)*(c + d).
Text Content:
Regular expressions are sequences of characters that define search patterns, primarily used for pattern matching within strings. The fundamental operators in regular expressions include concatenation (placing expressions sequentially), alternation (denoted by ‘—‘ or ‘+‘, representing a choice between expressions), and the Kleene star (denoted by ‘*‘, indicating zero or more repetitions of the preceding element). Understanding these operators is crucial for constructing expressions that accurately represent specific string patterns. For instance, the expression ‘(a—b)*‘ denotes any combination of the characters ’a’ and ’b’, including the empty string, while ‘a*‘ matches any number of consecutive ’a’ characters, including none. Mastery of these operators enables the creation of complex patterns for efficient text processing and matching tasks.
Example Output1:
    "Question": "When trying to construct a
    regular expression to match strings that
    start with an optional single occurrence
    of either ’e’ or ’f’, followed by zero
    or more repetitions of a group containing
    either ’g’ or ’h’ and then either ’i’ or
    ’j’, and end with exactly three
    repetitions of either ’k’ or ’l’, which
    of the following options would be the
    correct regular expression? ",
    "A": "(e|f)?((g|h)*(i|j))(k|l)^3",
    "B": "((e|f)+(g|h))*(i|j)(k|l)^3",
    "C": "(e|f)(g|h)*(i|j)^3(k|l)^3",
    "D": "(e|f)*((g|h)+(i|j))(k|l)^3",
    "Answer": "A"
[…]
current question:
[…]
Knowledge Point:
Register allocation and usage in assembly language
Main Idea: Determining the minimum number of clock cycles required to compute the expression $AB+ABC+BC$ on a pipelined RISC machine, considering operation latencies, data dependencies, and constraints on register modification.
Text Content:
In assembly language programming, efficient register allocation and usage are crucial for optimizing performance, especially in pipelined RISC architectures. Registers are limited, high-speed storage locations within the CPU that hold data for immediate processing, reducing the need to access slower main memory. Effective register allocation involves assigning these registers to frequently used variables and intermediate results, minimizing memory access and enhancing execution speed. This process can be complex due to the limited number of registers and the need to manage data dependencies and instruction latencies. Techniques such as graph coloring are often employed to allocate registers efficiently by modeling interference between variables and ensuring that no two simultaneously live variables share the same register. Additionally, understanding the specific roles and constraints of registers in a given architecture is essential, as certain instructions may require operands to reside in specific registers. By carefully managing register allocation and usage, assembly language programmers can significantly improve the efficiency and performance of their code.
Example Output2:
    "Question": "When computing the
    expression ( AB + ABC + BC ) on the
    described pipelined RISC machine, if
    we consider not only the operation
    latencies and register modification
    constraints but also the impact of
    register allocation strategies like
    graph coloring on reducing data
    dependencies, which of the following
    operation sequences would be the most
    efficient in terms of minimizing the
    number of clock cycles while also
    ensuring the best utilization of
    registers considering potential
    interference between variables according
    to the principles of register allocation
    in assembly language, and what is that
    minimum number of cycles?",
    "A": "Perform MUL R0, R1, R3; then use
    graph coloring to reallocate R1 for
    another operation to avoid interference;
    MUL R3, R2, R4; ADD R3, R2, R5; ADD R4,
    R5, R6; 5 cycles",
    "B": "First apply graph coloring to
    allocate registers optimally for the
    initial MUL operations; perform MUL
    R0, R1, R3; MUL R1,
    R2, R4; then handle ADD operations
    considering data dependencies based on
    the coloring result; ADD R3, R4, R5;
    ADD R5, R2, R6; 6 cycles",
    "C": "Without using graph coloring, just
    follow a traditional register allocation
    based on operand order; perform MUL R0, R1,
    R3; MUL R3, R2, R4; ADD R3, R2, R5; ADD R4,
    R5, R6; 7 cycles",
    "D": "Use graph coloring but in a
    suboptimal way that leads to some
    unnecessary register swaps; perform MUL
    R0, R1, R3; then swap R1 with another
    register due to the coloring plan; MUL
    R1, R2, R4; ADD R3, R4, R5; ADD
    R5, R2, R6; 8 cycles",
    "Answer": "B"
Finally, remember not to output in the following format (do not begin with “‘json and end with “‘).
‘‘‘json
    "Question": "When trying to construct a
    regular expression to match strings that
    start with an optional single occurrence
    of either ’e’ or ’f’, followed by zero or
    more repetitions of a group containing
    either ’g’ or ’h’ and then either ’i’ or
    ’j’, and end with exactly three repetitions
    of either ’k’ or ’l’, which of the
    following options would be the correct
    regular expression? ",
    "A": "(e|f)?((g|h)*(i|j))(k|l)^3",
    "B": "((e|f)+(g|h))*(i|j)(k|l)^3",
    "C": "(e|f)(g|h)*(i|j)^3(k|l)^3",
    "D": "(e|f)*((g|h)+(i|j))(k|l)^3",
    "Answer": "A"
Appendix C Question Design Example Based on Bloom’s Taxonomy
Layer: Remembering Question: Which symbol in regular expressions denotes zero or more repetitions of the preceding element? A: ? B: * C: + D: . Answer: B Layer: Understanding Question: In the context of regular expressions, what does the expression ‘(a—b)*‘ represent? A: Any string that contains only ’a’ repeated any number of times B: Any string that contains only ’b’ repeated any number of times C: Any combination of ’a’ and ’b’, including the empty string D: Any string that starts with ’a’ or ’b’ and ends with ’a’ or ’b’ Answer: C Layer: Applying Question: Given the regular expression ‘(a* + b)*(c + d)‘, which of the following regular expressions correctly applies the same pattern? A: (a—b)*(c—d) B: (a+b)*(c+d) C: (a*b)*(c+d) D: (a+b)*(c—d) Answer: D Layer: Analyzing Question: How does the regular expression ‘(a—b)*‘ differ from ‘a*‘ in terms of the possible strings it can match? A: They match exactly the same set of strings B: ‘(a—b)*‘ can match strings with ’b’, while ‘a*‘ cannot C: ‘a*‘ can match strings with ’b’, while ‘(a—b)*‘ cannot D: They both match strings with any number of ’a’s including ’b’s Answer: B
Layer: Evaluating Question: Considering the regular expression ‘(a* + b)*(c + d)‘, evaluate which of the following statements best justifies the use of the alternation operator ‘—‘ over the concatenation operator ‘+‘ in this context? A: The ‘—‘ operator allows for a choice between ’a’ and ’b’, which is necessary for matching a pattern that could start with either ’a’ or ’b’ B: The ‘+‘ operator is used for one or more repetitions, which is not suitable for optional elements C: The ‘—‘ operator is more efficient for matching a single character D: The ‘—‘ operator is used to match either ’a’ or ’b’, but it is not necessary for the pattern to be valid Answer: A Layer: Creating Question: If you were to create a new regular expression based on the pattern ‘(a* + b)*(c + d)‘ that matches a string starting with zero or more ’a’s, followed by one or more ’b’s, and ending with either ’c’ or ’d’, what would that regular expression be?, A: (a*b+)(c—d) B: (a+b*)(c—d) C: (a* + b+)(c—d) D: (a+b*)(c+d) Answer: C

### v3

@@
 To address this, we propose AdEval (Alignment-based Dynamic Evaluation), a dynamic data evaluation method aimed at mitigating the impact of data contamination on evaluation reliability.
-AdEval extracts key knowledge points and main ideas to align dynamically generated questions with static data’s core concepts.
-It also leverages online search to provide detailed explanations of related knowledge points, thereby creating high-quality evaluation samples with robust knowledge support.
-Furthermore, AdEval incorporates mechanisms to control the number and complexity of questions, enabling dynamic alignment and flexible adjustment.
-This ensures that the generated questions align with the complexity of static data while supporting varied complexity levels.
-Based on Bloom’s taxonomy, AdEval conducts a multi-dimensional evaluation of LLMs across six cognitive levels: remembering, understanding, applying, analyzing, evaluating, and creating.
 Experimental results on multiple datasets demonstrate that AdEval effectively reduces the impact of data contamination on evaluation outcomes, enhancing both the fairness and reliability of the evaluation process.
@@
 These methods explore the potential of dynamic evaluation from different perspectives, offering diverse tools for assessing the capabilities of LLMs.
-To address these challenges, this paper proposes AdEval (Alignment-based Dynamic Evaluation), a dynamic data evaluation method designed to reduce the impact of data contamination on LLMs evaluation results by dynamically generating aligned evaluation samples.
-The key innovations of AdEval are as follows:
-  1. 1.
-    Dynamically generating aligned evaluation data: AdEval extracts key knowledge points and main ideas from static data and combines them with online search to provide detailed expansions of related content.
-This process generates dynamic data that aligns with the core concepts of the static data, ensuring both dynamism and consistency.
-  2. 2.
-    Complexity control and dynamic alignment: AdEval incorporates a complexity regulation mechanism to dynamically adjust the difficulty of questions.
-This ensures that the generated data matches the difficulty level of static data while accommodating diverse complexity requirements.
-  3. 3.
-    Multi-dimensional cognitive evaluation: Using Bloom’s taxonomy, AdEval evaluates LLMs across six cognitive dimensions—remembering, understanding, applying, analyzing, evaluating, and creating—offering a comprehensive assessment of their performance in various cognitive tasks.
-Experimental results demonstrate that AdEval effectively reduces the impact of data contamination on evaluation outcomes, showcasing significant advantages in the reliability, fairness, and diversity of the generated data.
-Figure 1: An Overview of the AdEval Process: Generating Dynamic Datasets from Static Data.
 II Related Work
@@
 III Methods
-The proposed process for dynamically generating datasets is shown in Figure 1 , consisting of the following steps: data preprocessing, extraction of knowledge points and main ideas, online search, question design,and complexity control , quality checking.
+The proposed process for dynamically generating datasets is shown in , consisting of the following steps: data preprocessing, extraction of knowledge points and main ideas, online search, question design,and complexity control , quality checking.
 III-A Prompt Optimization
@@
 The floor function ensures valid integer indices.
-III-C Knowledge Point and Main Idea Extraction
-In this section, the relevant knowledge points are extracted from the sampled dataset.
-During the extraction process, LLMs autonomously generate knowledge points, which need to meet the following requirements:
-  1. 1.
-    The generated knowledge points must be relevant to the question.
-  2. 2.
-    They should focus on summarizing broad concepts or categories of knowledge related to the problem rather than detailed explanations.
-  3. 3.
-    The generated content must adhere to a predefined format. If it does not, the LLM is re-invoked to generate compliant results.
-Following the extraction of knowledge points, the main idea of the question is summarized. The LLM generates the main idea by summarizing the question, adhering to the following two requirements:
-  1. 1.
-    The main idea must encompass the context of the problem and the core of the answer.
-  2. 2.
-    The content must be concise and clear, avoiding redundant expressions.
-The main idea plays a key role in the subsequent online search and question design stages, ensuring that the newly generated questions align with the core ideas of the static dataset questions.
-During the extraction of knowledge points and main ideas, a few-shot strategy is employed, providing examples tailored to different datasets to improve the quality of the LLM-generated content.
-Overall, the knowledge points depict various aspects of the current question, while the main idea focuses on its core concept.
-This design ensures consistency in the core ideas while expanding the content of new questions in different directions.
-The prompts and few-shot examples for extracting knowledge points and main ideas are detailed in Appendices  A-A,  A-B and  B-A,  B-B.
-Below is an example of knowledge point and main idea extraction.
-Extract Knowledge Point and Main Idea Example Input Question: Which of the following regular expressions is equivalent to (describes the same set of strings as) (a* + b)*(c + d)?
-?
-A a*(c + d)+ b(c + d).
-B a*(c + d)* + b(c + d)*.
-C a*(c + d)+ b*(c + d).
-D (a + b)*c +(a + b)*d.
-Answer: D
-Example Output Knowledge Point: [”Regular expressions and their operators”, ”Concatenation in regular expressions”, ”Union operator (+) in regular expressions”, ”Kleene star (*) in regular expressions”, ”Equivalence of regular expressions”, ”Pattern matching with regular expressions”] Main Idea: Finding the regular expression equivalent to (a* + b)*(c + d).
-III-D Online Search for Knowledge Point Explanations
-In this section, we utilize the online search function of large language models to input the problem statement, related knowledge points, and corresponding main ideas of each question in the static dataset.
-The model then generates detailed explanations for the knowledge points.
-The generated content needs to meet the following three requirements:
-  1. 1.
-    Focus on the Knowledge Points: The detailed explanation must be closely centered around the knowledge points provided, and can include background or supplementary information from the question or main ideas, but the focus must remain on the knowledge points.
-  2. 2.
-    Concise Paragraphs: The output must be a clear and coherent paragraph, avoiding overly long or excessively short content, and must not include paragraph breaks or excessive embellishments.
-  3. 3.
-    Straightforward: The output must be comprehensive, accurate, and logically clear, focusing on the main topic while avoiding irrelevant information or vague statements.
-Each question is associated with a main idea and multiple knowledge points.
-To increase the flexibility of generation, we designed a knowledge point selection mechanism that controls the number of knowledge points selected.
-In practical tests, we found that generating a detailed explanation for one knowledge point using the Qwen-Plus API takes about 30 seconds.
-If explanations are generated for each knowledge point, it would lead to considerable time and cost consumption.
-To address this, we introduced a knowledge point selection strategy: we randomly select a specified number of knowledge points from the array of knowledge points for processing.
-For example, if the parameter is set to 2, and the number of knowledge points exceeds 2, 2 points are selected at random; if there are fewer than 2, all knowledge points are selected.
-We then traverse the selected knowledge points and generate related explanations based on the problem statement and main ideas.
-The online search prompts and few-shot examples are detailed in Appendix  A-C and  B-C. Below is an example of the online search for knowledge point explanations.
-Online Search Example Example Output Regular expressions are sequences of characters that define search patterns, primarily used for pattern matching within strings.
-The fundamental operators in regular expressions include concatenation (placing expressions sequentially), alternation (denoted by ‘—‘ or ‘+‘, representing a choice between expressions), and the Kleene star (denoted by ‘*‘, indicating zero or more repetitions of the preceding element).
-Understanding these operators is crucial for constructing expressions that accurately represent specific …
-III-E Question Design
-This section aims to generate two types of new questions by embedding the aforementioned static dataset questions, knowledge points, main ideas, and knowledge point explanations into the prompts: general new questions and Bloom’s Taxonomy-based[41, 42] new questions.
-The generated questions needs to meet the following requirements:
-  1. 1.
-    High Relevance: The question, correct option, and answer must be directly based on the provided text and closely related to the main idea.
-  2. 2.
-    Consistent/Greater/Lesser Difficulty: The complexity of the question should match the reference question’s complexity, be slightly higher, or slightly lower, ensuring that the question requires deep understanding, analysis, and synthesis of details to arrive at the correct answer, rather than being a simple recall of facts.
-  3. 3.
-    Reasonable and Discriminative Options: The options should be logically clear, avoiding obvious answers.
-At least one strong distractor should be included to assess the student’s attention to detail and reasoning ability.
-  4. 4.
-    Avoid External Dependencies: The question and options must be entirely based on the provided text and main idea.
-If external content is required, it should be embedded directly into the question, avoiding references such as ”According to the text” or similar phrases.
-Questions should not rely on context that makes them difficult to understand in isolation.
-  5. 5.
-    Unique Answer: Ensure only one option is correct, with other options being plausible but subtly incorrect or only partially matching the main idea and text content.
-  6. 6.
-    Assess Deep Understanding: The designed questions can involve inference, comparison, method analysis, application of key concepts, etc., to increase the difficulty.
-  7. 7.
-    Innovative Question Stem: The question format should not be too similar to the reference questions, demonstrating your ability to innovate, while maintaining rigor and professionalism in the question style.
-  8. 8.
-    Different Emphasis: The generated questions should focus on different aspects compared to the reference questions.
-For example, if the reference question focuses on concept definitions, the generated question could focus on application of methods, comparison of key details, analysis of pros and cons, or practical scenario judgment from other angles.
-  9. 9.
-    Focus on Knowledge Points: The question should be centered around the provided knowledge points, with the core content tightly aligned with the main idea.
-The newly generated questions based on Bloom’s taxonomy, in addition to meeting the requirements 1-5 and 9 listed above, must also satisfy the following: Level Coverage: Each question corresponds to a cognitive level, covering all six levels of Bloom’s taxonomy (for specific details, see Appendix  A-E).
-There should be one question for each level, totaling six questions.
-The question design prompts and few-shot examples are detailed in Appendices  A-D,  A-E, and  B-D,  B-E.
-Below is an example of a new question based on Bloom’s Taxonomy(The complete example can be found in Appendix  C):
-Question Design Example Based on Bloom’s Taxonomy Layer: Remembering Question: Which symbol in regular expressions denotes zero or more repetitions of the preceding element?
-A: ?
-B: * C: + D: .
-Answer: B Layer: Creating Question: If you were to create a new regular expression based on the pattern ‘(a* + b)*(c + d)‘ that matches a string starting with zero or more ’a’s, followed by one or more ’b’s, and ending with either ’c’ or ’d’, what would that regular expression be?, A: (a*b+)(c—d) B: (a+b*)(c—d) C: (a* + b+)(c—d) D: (a+b*)(c+d) Answer: C
-III-F Question Complexity Control
-If no constraints are applied to the LLM’s question generation process, the results tend to produce relatively simple datasets.
-To solve this issue, we designed a question difficulty adjustment process (as shown in Figure 3) to ensure that the generated dataset’s complexity aligns with that of the static dataset.
-Figure 3: Question Difficulty Adjustment Process - Aligning with Static Dataset Example
-In the initial phase of generating the dataset, we introduce a complexity alignment mechanism through few-shot prompting to ensure the complexity of the generated data matches expectations.
-However, in the initial phase, there is a large discrepancy in complexity, so further adjustments are necessary to control the complexity of the dataset within an acceptable error range.
-Specifically, we indirectly measure the complexity of the dataset using the accuracy of the preset model (e.g., Qwen-7B-Chat).
-The newly generated dataset is tested with the preset model, and the accuracy of the model is calculated to derive the deviation in question difficulty:
-$\Delta=s_{\text{target}}-s_{\text{gen}}$ (2)
-Where $s_{\text{target}}$ represents the desired accuracy (for example, when aligning with the static dataset, it refers to the accuracy of the static dataset on the preset model), and $s_{\text{gen}}$ represents the accuracy of the newly generated dataset on the preset model.
-We aim for the deviation to satisfy $|\Delta|<5\%$.
-Adjustment Strategy:
-If $\Delta>5\%$: The generated dataset’s complexity is too high.
-We then select the questions that the model answered incorrectly from the generated dataset and reconstruct them to simplify the questions.
-If $\Delta<-5\%$: The generated dataset’s complexity is too low. We then select the questions that the model answered correctly and reconstruct them to increase the complexity.
-After each reconstruction, the newly generated dataset is re-tested, and the above steps are repeated until $|\Delta|<5\%$.
-Through this process, we can effectively control the complexity of the generated dataset, aligning it with the static dataset or preset target.
-The complexity control prompts and few-shot examples are detailed in Appendices  A-F and  B-F.
-III-G Quality Control
-To ensure that each step of the AdEval process meets quality requirements, we introduced a multi-model voting mechanism for quality assessment.
-Specifically, multiple LLMs (including Doubao-pro-32k, Qwen-plus, and GPT-4o-mini) are used for quality evaluation.
-The voting rules are as follows:
-  1. 1.
-    If at least 2 out of the 3 responses are judged as ”pass,” the question is considered of acceptable quality.
-  2. 2.
-    If the above condition is not met, the question is deemed to have quality issues and is regenerated.
-The quality control prompts are detailed in Appendix  A-G.
 IV Experiment